\begin{thebibliography}{10}

\bibitem{ashikhmin1998minimal}
A~Ashikhmin and Alexander Barg.
\newblock Minimal vectors in linear codes.
\newblock {\em IEEE Transactions on Information Theory}, 44(5):2010--2017,
  1998.

\bibitem{calderbank1984three}
AR~Calderbank and JM~Goethals.
\newblock Three-weight codes and association schemes.
\newblock {\em Philips Journal of Research}, 39(4-5):143--152, 1984.

\bibitem{calderbank1986geometry}
R~Calderbank and WM~Kantor.
\newblock The geometry of two-weight codes.
\newblock {\em Bulletin of the London Mathematical Society}, 18(2):97--122,
  1986.

\bibitem{carlet2005linear}
Claude Carlet, Cunsheng Ding, and Jin Yuan.
\newblock Linear codes from perfect nonlinear mappings and their secret sharing
  schemes.
\newblock {\em IEEE Transactions on Information Theory}, 51(6):2089--2102,
  2005.

\bibitem{ding2008two}
Cunsheng Ding, Jinquan Luo, and Harald Niederreiter.
\newblock Two-weight codes punctured from irreducible cyclic codes.
\newblock In {\em Coding And Cryptology}, pages 119--124. World Scientific,
  2008.

\bibitem{ding2007cyclotomic}
Cunsheng Ding and Harald Niederreiter.
\newblock Cyclotomic linear codes of order 3.
\newblock {\em IEEE Transactions on Information Theory}, 53(6):2274--2277,
  2007.

\bibitem{ding2005coding}
Cunsheng Ding and Xuesong Wang.
\newblock A coding theory construction of new systematic authentication codes.
\newblock {\em Theoretical Computer Science}, 330(1):81--99, 2005.

\bibitem{ding2003covering}
Cunsheng Ding and Jin Yuan.
\newblock Covering and secret sharing with linear codes.
\newblock In {\em Discrete Mathematics and Theoretical Computer Science}, pages
  11--25. Springer, 2003.

\bibitem{ding2014binary}
Kelan Ding and Cunsheng Ding.
\newblock Binary linear codes with three weights.
\newblock {\em IEEE Communications Letters}, 18(11):1879--1882, 2014.

\bibitem{ding2015class}
Kelan Ding and Cunsheng Ding.
\newblock A class of two-weight and three-weight codes and their applications
  in secret sharing.
\newblock {\em IEEE Transactions on Information Theory}, 61(11):5835--5842,
  2015.

\bibitem{griesmer1960bound}
James~H Griesmer.
\newblock A bound for error-correcting codes.
\newblock {\em IBM Journal of Research and Development}, 4(5):532--542, 1960.

\bibitem{heng2018construction}
Ziling Heng and Cunsheng Ding.
\newblock A construction of $q$-ary linear codes with irreducible cyclic codes.
\newblock {\em Designs, Codes and Cryptography}, pages 1--22, 2018.

\bibitem{heng2015class}
Ziling Heng and Qin Yue.
\newblock A class of binary linear codes with at most three weights.
\newblock {\em IEEE Communications Letters}, 19(9):1488--1491, 2015.

\bibitem{heng2016two}
Ziling Heng and Qin Yue.
\newblock Two classes of two-weight linear codes.
\newblock {\em Finite Fields and Their Applications}, 38:72--92, 2016.

\bibitem{heng2017construction}
Ziling Heng and Qin Yue.
\newblock A construction of $q$-ary linear codes with two weights.
\newblock {\em Finite Fields and Their Applications}, 48:20--42, 2017.

\bibitem{heng2016three}
Ziling Heng, Qin Yue, and Chengju Li.
\newblock Three classes of linear codes with two or three weights.
\newblock {\em Discrete Mathematics}, 339(11):2832--2847, 2016.

\bibitem{lidl1997finite}
Rudolf Lidl and Harald Niederreiter.
\newblock {\em Finite fields}, volume~20.
\newblock Cambridge university press, 1997.

\bibitem{liu2017several}
Hongwei Liu and Youcef Maouche.
\newblock Several classes of trace codes with either optimal two weights or a
  few weights over $\mathbb{F}_q+ u\mathbb{F}_q$.
\newblock {\em arXiv preprint arXiv:1703.04968}, 2017.

\bibitem{liu2018several}
Hongwei Liu and Youcef Maouche.
\newblock Several new classes of linear codes with few weights.
\newblock {\em Cryptography and Communications}, pages 1--10, 2018.

\bibitem{liu2018two}
Yan Liu, Minjia Shi, and Patrick Sol{\'e}.
\newblock Two-weight and three-weight codes from trace codes over
  $\mathbb{F}_p+u \mathbb{F}_p+v \mathbb{F}_p+uv \mathbb{F}_p$.
\newblock {\em Discrete Mathematics}, 341(2):350--357, 2018.

\bibitem{luo2018five}
Gaojun Luo and Xiwang Cao.
\newblock Five classes of optimal two-weight linear codes.
\newblock {\em Cryptography and Communications}, 10(6):1119--1135, 2018.

\bibitem{luo2018binary}
Gaojun Luo, Xiwang Cao, Shanding Xu, and Jiafu Mi.
\newblock Binary linear codes with two or three weights from niho exponents.
\newblock {\em Cryptography and Communications}, 10(2):301--318, 2018.

\bibitem{myerson1981period}
Gerald Myerson.
\newblock Period polynomials and {Gauss} sums for finite fields.
\newblock {\em Acta Arithmetica}, 39(3):251--264, 1981.

\bibitem{shi2016optimal}
Minjia Shi, Yan Liu, and Patrick Sol{\'e}.
\newblock Optimal two-weight codes from trace codes over
  $\mathbb{F}_2+u\mathbb{F}_2$.
\newblock {\em IEEE Communications Letters}, 20(12):2346--2349, 2016.

\bibitem{shi2017two}
Minjia Shi, Rongsheng Wu, Yan Liu, and Patrick Sol{\'e}.
\newblock Two and three weight codes over $\mathbb{F}_p+ u \mathbb{F}_p$.
\newblock {\em Cryptography and Communications}, 9(5):637--646, 2017.

\bibitem{shi2016new}
Minjia Shi, Rongsheng Wu, Liqin Qian, Lin Sok, and Patrick Sol{\'e}.
\newblock New classes of $p$-ary few weight codes.
\newblock {\em Bulletin of the Malaysian Mathematical Sciences Society}, pages
  1--20, 2016.

\bibitem{tang2016linear}
Chunming Tang, Nian Li, Yanfeng Qi, Zhengchun Zhou, and Tor Helleseth.
\newblock Linear codes with two or three weights from weakly regular bent
  functions.
\newblock {\em IEEE Transactions on Information Theory}, 62(3):1166--1176,
  2016.

\bibitem{torleiv2007codes}
Klove Torleiv.
\newblock {\em Codes For Error Detection}, volume~2.
\newblock World Scientific, 2007.

\bibitem{yuan2006secret}
Jin Yuan and Cunsheng Ding.
\newblock Secret sharing schemes from three classes of linear codes.
\newblock {\em IEEE Transactions on Information Theory}, 52(1):206--212, 2006.

\bibitem{zhou2015linear}
Zhengchun Zhou, Nian Li, Cuiling Fan, and Tor Helleseth.
\newblock Linear codes with two or three weights from quadratic bent functions.
\newblock {\em Designs, Codes and Cryptography}, pages 1--13, 2015.

\end{thebibliography}

\documentclass{svjour3}

\usepackage{amsmath,amssymb}
\usepackage{hyperref}
\usepackage{subfig}

\usepackage[amsmath, thmmarks]{ntheorem}
{\theoremstyle{nonumberplain}
\theoremsymbol{\ensuremath{\Box}}
\newtheorem{pf}{Proof}}

\newtheorem{thm}{Theorem}
\newtheorem{lem}{Lemma}
\newtheorem{cll}{Corollary}

\newtheorem{ex}{Example}

\title{A class of linear codes with few weights}

\author{Gaopeng Jian}    

\institute{Gaopeng Jian \at Key Laboratory of Machine Perception(MOE), School of EECS, Peking University, Beijing 100871, P.R.China; \\
\email{gpjian@pku.edu.cn}
}

\date{}

\begin{document}
%
%
%
  
  

\maketitle  

\begin{abstract}
Linear codes have been an interesting topic in both theory and practice for many years. In this paper, a class of $q$-ary linear codes with few weights are presented and their weight distributions are determined using Gauss periods.
Some of the linear codes obtained are optimal or almost optimal  with respect to the Griesmer bound.
As s applications, these linear codes can be used to construct secret sharing schemes with nice access structures.
\keywords{Linear codes \and Weight distribution \and Gauss periods}
\end{abstract}



\section{Introduction}
Let $q$ be a power of a prime and $\mathbb{F}_q$ denote the finite field with $q$ elements.
An $[n,k,d]$ linear code $C$ over $\mathbb{F}_q$ is a $k$-dimensional subspace of $\mathbb{F}_q^n$ with minimum Hamming distance $d$.
Let $A_i$ be the number of codewords with Hamming weight $i$ in $C$. 
The polynomial $1+A_1 z+A_2 z^2+\cdots+A_n z^n$ is called the weight enumerator of $C$ and the sequence $(1, A_1, A_2, \ldots, A_n)$ called the weight distribution of $C$.
If the number of nonzero $A_i$ in the sequence $(A_1, A_2, \ldots, A_n)$ is equal to $t$, we call $C$ a $t$-weight code.
The weight distribution of a code contains important information on its error correcting capability and the error probability of its error detection and correction with respect to some algorithms \cite{torleiv2007codes}.
In addition, much attention has been paid to two-weight and three-weight linear codes  \cite{tang2016linear,zhou2015linear,heng2016three,heng2016two,heng2015class,heng2017construction,luo2018binary,ding2014binary,ding2015class} due to their applications in secret sharing schemes \cite{yuan2006secret,carlet2005linear}, strongly regular graphs \cite{calderbank1986geometry}, association schemes \cite{calderbank1984three} and authentication codes \cite{ding2005coding}.  

We recall the Griesmer bound for linear codes in the following lemma \cite{griesmer1960bound}.
\begin{lem}[Griesmer bound]  
For any $[n,k,d]$ code over $\mathbb{F}_q$, we have
\[
n \ge \sum_{i=0}^{k-1} \left \lceil \frac{d}{q^i} \right \rceil,
\]
where $\lceil \cdot \rceil$  is the ceiling function.
\end{lem}

Let $\text{Tr}^m_1(\cdot)$ denote the trace function from $\mathbb{F}_{q^m}$ to $\mathbb{F}_q$, and let $D=\{ d_1, d_2, \ldots, d_n \}$ be a subset of $\mathbb{F}_{q^m}^*$. 
Define a linear code of length $n = |D|$ over $\mathbb{F}_q$ by
\begin{equation} \label{d1}
C_D=\{(\text{Tr}^m_1(ad_1), \text{Tr}^m_1(ad_2), \ldots, \text{Tr}^m_1(ad_n)): a \in \mathbb{F}_{q^m} \}.
\end{equation}
The set $D$ is called the \emph{defining set} of the trace code $C_D$.
The construction was first proposed by Ding et al. in  \cite{ding2007cyclotomic} and many classes of known codes could be produced by properly selecting the defining set.
Recently, researchers have extended the construction to codes over finite rings and obtain many classes of linear codes with few weights by the Gray map \cite{luo2018five,liu2017several,shi2016optimal,shi2017two,shi2016new,liu2018two}.

Let $m_1$ be a positive divisor of $m$, and let $\text{Tr}^{m_1}_1(\cdot)$ denote the trace function from $\mathbb{F}_{q^{m_1}}$ to $\mathbb{F}_q$.
For $K \subset \mathbb{F}_{q^{m_1}} \times \mathbb{F}_{q^m}^*$, Liu et al. \cite{liu2018several} defined a linear code of length $2|K|$ over $\mathbb{F}_q$ by
\begin{equation} \label{e1}
C_K=\{c(a,b): a \in \mathbb{F}_{q^{m_1}}, b \in \mathbb{F}_{q^m} \}, 
\end{equation}
where
\[
c(a,b)=(\text{Tr}^{m_1}_1(ax)+\text{Tr}^m_1(by)|\text{Tr}^{m_1}_1(ax)+\text{Tr}^m_1((a+b)y))_{(x,y) \in K}.
\]
Note that the Gray image of trace codes over the ring $R=\mathbb{F}_q+u\mathbb{F}_q$ with $u^2=0$ coincides with the construction when $m=m_1$.

Let $\omega$ be a fixed primitive element of $\mathbb{F}_{q^m}$, and let $D=\{1,\omega^h,\ldots,\omega^{(n-1)h} \}$, where $h | \frac{q^m-1}{q-1}$ and $\frac{q^m-1}{h(q-1)} |n$.
Note that if $n=\frac{q^m-1}{h}$, then $D=C_0^{(h,q^m)}$ and the linear code $C_D$ defined by \eqref{d1} is equivalent to an irreducible cyclic codes over $\mathbb{F}_q$, which is not true in general \cite{ding2008two,heng2018construction}.
In this paper we will investigate the weight distribution of $C_K$ defined by \eqref{e1} with $K=\mathbb{F}_{q^{m_1}} \times D$.
Several classes of two-weight and three-weight linear codes are derived employing Gauss periods over finite fields.
In particular, the two-weight codes are optimal with respect to the Griesmer bound if $n$ is small.
Furthermore, an application to secret sharing schemes is sketched out.

This paper is organized as follows. 
In Section \ref{pl}, we briefly recall some definitions and notations which will be used later.  
In Section \ref{wt}, we investigate the weight distribution of $C_K$ by using Gauss periods and construct two-weight and three-weight codes over $\mathbb{F}_q$. 
In Section \ref{sss}, we state an application in secret sharing schemes.
In Section \ref{clu}, we conclude this paper.

\section{Preliminaries} \label{pl}
In this section,  we recall some basic results of characters, cyclotomic classes and Gauss periods. Interested readers are referred to \cite{lidl1997finite,myerson1981period} for more details.

Suppose that $q=p^s$ for an odd prime $p$ and a positive integer $s$. For $a \in \mathbb{F}_q$, an additive character of the finite field $\mathbb{F}_q$ can be defined by
\[
\chi_a: \mathbb{F}_q \rightarrow \mathbb{C}^*, \chi_a(x)=\zeta_p^{\text{Tr}(ax)},
\]
where $\zeta_p=e^{\frac{2 \pi \sqrt{-1}}{p}}$ is a primitive $p$-th root of unity and $\text{Tr}(\cdot)$ denotes the trace function from $\mathbb{F}_q$ onto $\mathbb{F}_p$. 
It is clear that $\chi_0(x) = 1$ for all $x \in \mathbb{F}_q$ and $\chi_0$ is called the \emph{trivial additive character} of $\mathbb{F}_q$. If $a=1$, we call $\chi:=\chi_1$ the \emph{canonical additive character} of $\mathbb{F}_q$. 
The orthogonal property of additive characters is given as
\[
\sum_{x \in \mathbb{F}_q} \chi(ax)=\begin{cases}
q, & \text{if } a=0, \\
0, & \text{otherwise}.
\end{cases}
\] 

Let $\alpha$ be a fixed primitive element of $\mathbb{F}_q$. For a positive integer $N>1$ such that $N|(q-1)$, we define
\[
C_i^{(N,q)}=\alpha^i \langle \alpha^N \rangle, \ i=0,1,\ldots,N-1,
\]
where $\langle \alpha^N \rangle$ denotes the cyclic subgroup of $\mathbb{F}_q^*$ generated by $\alpha^N$.
The cosets $C_i^{(N,q)}$ are called the \emph{cyclotomic classes} of order $N$ in  $\mathbb{F}_q$. 

The \emph{Gaussian periods} are defined by
\[
\eta_i^{(N,q)}=\sum_{x \in C_i^{(N,q)}} \chi(x),\ i=0,1,\ldots,N-1,
\]
where $\chi$ is the canonical additive character of $\mathbb{F}_q$.

\begin{lem} \label{l3}
\[
\left | \eta_i^{(N,q)}+\frac{1}{N} \right | \le \frac{(N-1)\sqrt{q}}{N}. 
\]
\end{lem}

\begin{lem}  \label{l2}
Let $N=2$, then the Gaussian periods are given by
\[
\eta_0^{(2,q)}=\begin{cases}
\frac{-1+(-1)^{s-1}\sqrt{q}}{2}, & \text{if} \ p \equiv 1 \pmod{4},\\
\frac{-1+(-1)^{s-1}(\sqrt{-1})^{s} \sqrt{q}}{2}, & \text{if} \ p \equiv 3 \pmod{4}
\end{cases}
\]
and
\[
\eta_1^{(2,q)}=-1-\eta_0^{(2,q)}.
\]
\end{lem}

\begin{lem}[The semiprimitive case] \label{l7}
Assume that $N \ge 2$, $q=p^{2j\gamma}$, where $N | (p^j+1)$ and $j$ is the smallest such positive integer. Then the Gaussian periods of order $N$ are given below:
\begin{description}
\item[(a)] If $\gamma, p, \frac{p^j+1}{N}$ are all odd, then
\[
\eta_{\frac{N}{2}}^{(N,q)}=\sqrt{q}-\frac{\sqrt{q}+1}{N}, \
\eta_i^{(N,q)}=-\frac{\sqrt{q}+1}{N}  \text{ for all } i \neq \frac{N}{2}.
\]
\item[(b)] In all the other cases,
\[
\eta_0^{(N,q)}=\frac{(-1)^{\gamma+1}(N-1)\sqrt{q}-1}{N}, \
\eta_i^{(N,q)}=\frac{(-1)^{\gamma}\sqrt{q}-1}{N} \text{ for all } i \neq 0.
\]
\end{description}
\end{lem}

\section{The weight distribution of  $C_K$} \label{wt}
In this section, we investigate the weight distribution of $C_K$ defined by \eqref{e1}.
Let $\mathbb{F}_{q^m}^*=\langle \omega \rangle$ and $D=\{\omega^{hi}:0 \le i<n \}$ where $h | \frac{q^m-1}{q-1}$ and $\frac{q^m-1}{h(q-1)}|n$.
We know that any codeword of $C_K$ has the following form:
\[
c(a,b)=(\text{Tr}^{m_1}_1(ax)+\text{Tr}^m_1(by)|\text{Tr}^{m_1}_1(ax)+\text{Tr}^m_1((a+b)y))_{(x,y) \in K},
\]
where $K=\mathbb{F}_{q^{m_1}} \times D$ and $a \in \mathbb{F}_{q^{m_1}}$, $b \in \mathbb{F}_{q^m}$.
The Hamming weight of $c(a,b)$ is expressed in terms of  Gauss periods below.
\begin{thm} \label{th}
For $a \in \mathbb{F}_{q^{m_1}}$, $b \in \mathbb{F}_{q^m}$, the Hamming weight $w_H(c(a,b))$ is 
\begin{enumerate}
\item If $a=b=0$, then $w_H(c(a,b))=0$;
\item If $a=0$ and $b \in C_i^{(h,q^m)}$, $i=0,1,\ldots,h-1$, then 
\[
w_H(c(a,b))=\frac{2nq^{m_1-1}(q-1)}{q^m-1} \left( q^m-1-h\eta_i^{(h,q^m)} \right);
\]
\item If $a \in \mathbb{F}_{q^{m_1}}^*$ and $b \in \mathbb{F}_{q^m}$, then $w_H(c(a,b))=2nq^{m_1-1}(q-1)$.
\end{enumerate}
\end{thm}
\begin{pf}
Let $\chi$, $\chi^{(m)}$, $\chi^{(m_1)}$ be the canonical additive characters of $\mathbb{F}_q$, $\mathbb{F}_{q^m}$ and $\mathbb{F}_{q^{m_1}}$ respectively.
Let $N$ be a positive integer. For any vector $\mathbf{y}=(y_1,y_2,\ldots,y_N) \in \mathbb{F}_q^N$, let
\[
\Psi(\mathbf{y})=\sum_{i=1}^N \chi(y_i).
\]
By the orthogonal property of additive characters
\begin{align}
w_H(\mathbf{y}) &=N-\frac{1}{q}\sum_{i=1}^N \sum_{s \in \mathbb{F}_q} \chi(sy_i) \notag \\
&=(1-\frac{1}{q})N-\frac{1}{q} \sum_{s \in \mathbb{F}_q^*} \Psi(s\mathbf{y}) \label{wh}
\end{align}
If $a \in \mathbb{F}_{q^{m_1}}^*$ and $b \in \mathbb{F}_{q^m}$,
\begin{align}
\Psi(c(a,b))= &\sum_{x \in \mathbb{F}_{q^{m_1}}} \sum_{y \in D} \chi(\text{Tr}^{m_1}_1(ax)+\text{Tr}^m_1(by)) \notag \\
&+\sum_{x \in \mathbb{F}_{q^{m_1}}} \sum_{y \in D} \chi(\text{Tr}^{m_1}_1(ax)+\text{Tr}^m_1((a+b)y)) \notag\\
=&\sum_{y \in D}\chi^{(m)}(by)\sum_{x \in \mathbb{F}_{q^{m_1}}}\chi^{(m_1)}(ax) \notag\\
&+\sum_{y \in D}\chi^{(m)}((a+b)y)\sum_{x \in \mathbb{F}_{q^{m_1}}}\chi^{(m_1)}(ax) \notag \\
=&0. \label{w2}
\end{align}
If $a=0$, then $c(0,b) = (\text{Tr}^m_1(by)|\text{Tr}^m_1(by))_{(x,y) \in K}$ and
\begin{align*}
\sum_{s \in \mathbb{F}_q^*}\Psi(s \cdot c(0,b)) &=2\sum_{s \in \mathbb{F}_q^*}\sum_{x \in \mathbb{F}_{q^{m_1}}} \sum_{y \in D} \chi(s\text{Tr}^m_1(by)) \\
&=2q^{m_1}\sum_{s \in \mathbb{F}_q^*}\sum_{i=0}^{n-1}\chi^{(m)}(sb\omega^{hi}) 
\end{align*}
Since $h | \frac{q^m-1}{q-1}$ and $\frac{q^m-1}{h(q-1)}|n$, then 
\[
C_0^{(h,q^m)}=\langle \omega^h \rangle=\bigcup_{j=0}^{\frac{q^m-1}{h(q-1)}-1} \omega^{hj}\langle \omega^{\frac{q^m-1}{q-1}} \rangle=\bigcup_{j=0}^{\frac{q^m-1}{h(q-1)}-1} \omega^{hj} \mathbb{F}_q^*.
\]
and for $b \in C_i^{(h,q^m)}$,
\begin{align}
\sum_{s \in \mathbb{F}_q^*}\Psi(s \cdot c(0,b)) &=2q^{m_1} \sum_{s \in \mathbb{F}_q^*} \sum_{l=0}^{\frac{hn(q-1)}{q^m-1}-1} \sum_{j=0}^{\frac{q^m-1}{h(q-1)}-1} \chi^{(m)}(sb\omega^{hj+\frac{q^m-1}{q-1}l}) \notag \\
&=2q^{m_1} \sum_{l=0}^{\frac{hn(q-1)}{q^m-1}-1} \sum_{j=0}^{\frac{q^m-1}{h(q-1)}-1} \sum_{s \in \mathbb{F}_q^*} \chi^{(m)}((s\omega^{hj})b\omega^{\frac{q^m-1}{q-1}l}) \notag \\
&=2q^{m_1} \sum_{l=0}^{\frac{hn(q-1)}{q^m-1}-1} \sum_{t \in C_0^{(h,q^m)}} \chi^{(m)}(bt(\omega^h)^{\frac{q^m-1}{h(q-1)}l}) \notag \\
&=\frac{2q^{m_1}hn(q-1)}{q^m-1} \sum_{t \in C_0^{(h,q^m)}} \chi^{(m)}(bt) \notag \\
&=\frac{2q^{m_1}hn(q-1)}{q^m-1}\cdot \eta_i^{(h,q^m)}. \label{w1}
\end{align}
By \eqref{wh},\eqref{w2} and \eqref{w1} we complete the proof.
\end{pf}

\begin{cll}
If $h<q^{m/2}+1$, $C_K$ has  parameters
\[
\left[ 2nq^{m_1},m+m_1,d \ge \left \lceil \frac{2nq^{m/2+m_1-1}(q-1)}{q^m-1} (q^{m/2}+1-h) \right \rceil \right].
\]
\end{cll}
\begin{pf}
By Theorem \ref{th} and Lemma \ref{l3}, for $(a,b) \ne (0,0)$,
\[
w_H(c(a,b)) \ge \left \lceil \frac{2nq^{m_1-1}(q-1)}{q^m-1} (q^m-(h-1)q^{m/2}) \right \rceil>0
\]
as $h<q^{m/2}+1$.
Thus the dimension of $C_K$ is $m+m_1$.
\end{pf}

\begin{cll}\label{h1}
If $h=1$, $C_K$ is a $[2nq^{m_1},m+m_1,2nq^{m_1-1}(q-1)]$ two-weight linear code with the weight distribution given in Table \ref{t1}.
Furthermore, $C_K$ meets the Griesmer bound if $n<\frac{q(q^m-1)}{2(q-1)}$.
\end{cll}
\begin{pf}
If $h=1$, for $a=0$ and $b \in \mathbb{F}_{q^m}^*$,
\[
w_H(c(a,b))=\frac{2nq^{m_1-1}(q-1)}{q^m-1} \left( q^m-1-\eta_0^{(1,q^m)} \right)=\frac{2nq^{m_1+m-1}(q-1)}{q^m-1}.
\]
We can obtain the parameters and weight distribution of $C_K$ from Theorem \ref{th}.

If $n<\frac{q(q^m-1)}{2(q-1)}$, let $n=\frac{t(q^m-1)}{q-1}$ with $0<t<\frac{q}{2}$, then
\begin{align*}
\sum_{i=0}^{m+m_1-1} \left \lceil \frac{2nq^{m_1-1}(q-1)}{q^i} \right \rceil=& 2n(q-1)\sum_{i=0}^{m_1-1}q^{m_1-1-i}+\sum_{i=m_1}^{m+m_1-1} \left \lceil \frac{2t(q^m-1)}{q^{i+1-m_1}} \right \rceil \\
=& 2n(q^{m_1}-1)+\sum_{j=1}^m \left \lceil 2tq^{m-j}-\frac{2t}{q^j} \right \rceil \\
=&2n(q^{m_1}-1)+2t\sum_{j=1}^m q^{m-j} \\
=& 2nq^{m_1}.
\end{align*}
Thus $C_K$ meets the Griesmer bound.
\end{pf}

\begin{table}
\caption{Weight distribution for $h=1$} 
\centering
\begin{tabular}{ll} 
\hline
Weight                & Frequency \\
\hline
0 &1 \\
$2nq^{m_1-1}(q-1)$ & $q^{m_1+m}-q^m$\\
$\frac{2nq^{m_1+m-1}(q-1)}{q^m-1}$ & $q^m-1$\\
\hline
\end{tabular}
\label{t1}
\end{table}

\begin{ex}
Let $(q,m,m_1,h,n)=(3,2,1,1,4)$.
Then $C_K$ has parameters [24,3,16] and weight enumerator $1 + 18z^{16} + 8z^{18}$.
The code is optimal.
As a comparison, it is different from the best known linear codes from the Magma BKLC(GF(3),24,3) which has a different weight enumerator $1 + 12z^{16} + 12z^{17}+2z^{18}$.
\end{ex}

\begin{ex}
Let $(q,m,m_1,h,n)=(3,2,1,1,8)$.
Then $C_K$ has parameters [48,3,32] and weight enumerator $1 + 18z^{32} + 8z^{36}$.
The code is almost optimal as the best linear code of length 48 and dimension 3 over $\mathbb{F}_3$ has minimum weight 33.
\end{ex}

\begin{ex}
Let $(q,m,m_1,h,n)=(3,2,2,1,4)$.
Then $C_K$ has parameters [72,4,48] and weight enumerator $1 + 72z^{48} + 8z^{54}$.
The code is optimal.
As a comparison, it is different from the best known linear codes from the Magma BKLC(GF(3),72,4) which has a different weight enumerator $1 + 66z^{48} + 12z^{51}+2z^{54}$.
\end{ex}

\begin{cll}\label{h2}
If $h=2$, $C_K$ is a $[2nq^{m_1},m+m_1,\frac{2nq^{m/2+m_1-1}(q-1)}{q^m-1}(q^{m/2}-1)]$ three-weight linear code with the weight distribution given in Table \ref{t2}.
\end{cll}
\begin{pf}
Since $2|\frac{q^m-1}{q-1}$, $m$ is even.
For $a=0$ and $b \in C_i^{(2,q^m)}$, $i=0,1$, 
\[
w_H(c(a,b))=\frac{2nq^{m_1-1}(q-1)}{q^m-1}\left(q^m-(-1)^i(1+2\eta_0^{(2,q^m)})\right)
\]
and $|2\eta_0^{(2,q^m)}+1|=q^{m/2}$ by Lemma \ref{l2}.
We can obtain the parameters and weight distribution of $C_K$ from Theorem \ref{th}.
\end{pf}

\begin{table}
\caption{Weight distribution for $h=2$} 
\centering
\begin{tabular}{ll} 
\hline
Weight                & Frequency \\
\hline
0 &1 \\
$2nq^{m_1-1}(q-1)$ & $q^{m_1+m}-q^m$\\
$\frac{2nq^{m/2+m_1-1}(q-1)}{q^m-1}(q^{m/2}-1)$ & $\frac{q^m-1}{2}$\\
$\frac{2nq^{m/2+m_1-1}(q-1)}{q^m-1}(q^{m/2}+1)$ & $\frac{q^m-1}{2}$\\
\hline
\end{tabular}
\label{t2}
\end{table}

\begin{ex}
Let $(q,m,m_1,h,n)=(3,2,1,2,4)$.
Then $C_K$ has parameters [24,3,12] and weight enumerator $1 + 4z^{12} + 18z^{16} + 4z^{24}$.
\end{ex}

\begin{ex}
Let $(q,m,m_1,h,n)=(3,2,2,2,4)$.
Then $C_K$ has parameters [72,4,36] and weight enumerator $1 + 4z^{36} + 72z^{48} + 4z^{72}$.
\end{ex}

\begin{cll}\label{h3}
Assume that $2 \le h<q^{m/2}+1$, $q^m= p^{2j \gamma}$, where $h|(p^j +1)$ and $j$ is the smallest such positive integer. Then $C_K$ is a $[2nq^{m_1},m+m_1,d]$ three-weight linear code with the weight distribution given in Table \ref{t3}, where 
\[
d=\begin{cases} 
\frac{2nq^{m/2+m_1-1}(q-1)}{q^m-1}(q^{m/2}-1) & \text{if $\gamma$ is even} \\
\frac{2nq^{m/2+m_1-1}(q-1)}{q^m-1}(q^{m/2}-h+1) & \text{if $\gamma$ is odd.}
\end{cases}
\]
\end{cll}
\begin{pf}
By Lemma \ref{l7}, $\{\eta_i^{(h,q^m)}:i=0,1,\ldots,h-1\}$ takes value $\frac{(-1)^{\gamma+1}(h-1)q^{m/2}-1}{h}$ with multiplicity 1 and $\frac{(-1)^{\gamma}q^{m/2}-1}{h}$ with multiplicity $h-1$.
We can obtain the parameters and weight distribution of $C_K$ from Theorem \ref{th}.
\end{pf}

\begin{table}
\caption{Weight distribution in the semiprimitive case} 
\centering
\begin{tabular}{ll} 
\hline
Weight                & Frequency \\
\hline
0 &1 \\
$2nq^{m_1-1}(q-1)$ & $q^{m_1+m}-q^m$\\
$\frac{2nq^{m/2+m_1-1}(q-1)}{q^m-1}(q^{m/2}+(-1)^{\gamma}(h-1))$ & $\frac{q^m-1}{h}$\\
$\frac{2nq^{m/2+m_1-1}(q-1)}{q^m-1}(q^{m/2}+(-1)^{\gamma+1})$ & $\frac{(q^m-1)(h-1)}{h}$\\
\hline
\end{tabular}
\label{t3}
\end{table}

\begin{ex}
Let $(q,m,m_1,h,n)=(3,4,2,4,10)$. Then $j=1$ and $\gamma=2$. 
The code $C_K$ has parameters [180,6,108] and weight enumerator $1 + 60z^{108} + 648z^{120} + 20z^{162}$.
\end{ex}

\begin{ex}
Let $(q,m,m_1,h,n)=(3,4,2,5,8)$. Then $j=2$ and $\gamma=1$.
The code $C_K$ has parameters [144,6,54] and weight enumerator $1 + 16z^{54} + 648z^{96} + 64z^{108}$.
\end{ex}

\section{Applications in secret sharing schemes} \label{sss}
For a vector $\mathbf{x} \in \mathbb{F}_q^n$, the support $s(\mathbf{x})$ of $\mathbf{x}$ is defined as the set of indices where it is nonzero.
We say that a vector $\mathbf{x}$ covers a vector $\mathbf{y}$ if $s(\mathbf{x})$ contains $s(\mathbf{y})$.
For a linear code $C$ over $\mathbb{F}_q$, a codeword $\mathbf{c} \in C$ is minimal if it covers only codewords of the form $a \cdot \mathbf{c}$, where $a \in \mathbb{F}_q$.
$C$ is minimal if every codeword of $C$ is minimal.
If the weights of $C$ are close enough to each other, then $C$ is minimal, as described by the following lemma \cite{ashikhmin1998minimal}.

\begin{lem}
Denote by $w_{min}$ and $w_{max}$ the minimum and maximum nonzero weight of a given $q$-ary linear code $C$, respectively. If $w_{min}/w_{max}>\frac{q-1}{q}$, then $C$ is minimal.
\end{lem}

The notion of minimal codewords was introduced to determine the set of all minimal access sets of a secret sharing scheme (SSS).
Massey's scheme is a construction of a SSS based on coding theory \cite{yuan2006secret}.
When  $C$ is minimal, it was stated in \cite{ding2003covering} that there is the following alternative, depending on $d^{\perp}$ (the minimum distance of the dual of $C$):
\begin{itemize}
\item If $d^{\perp} \ge 3$, then the SSS is \emph{``democratic''}: every user belongs to the same number of coalitions,
\item If $d^{\perp}=2$, then there are users who belong to every coalition: the \emph{``dictators''}.
\end{itemize}
It's easy to check the following:
\begin{enumerate}
\item The code $C_K$ in Corollary \ref{h1} is minimal, provided that $m > 1$. 
\item The code $C_K$ in Corollary \ref{h2} is minimal, provided that $m > 2$.
\item The code $C_K$ in Corollary \ref{h3} is minimal, provided that $\gamma$ is even and $h(q-1)<q^{m/2}-1$ or $\gamma$ is odd and $hq <q^{m/2}+1$.
\end{enumerate}
Besides, a SSS built on $C_K$ is dictatorial as the minimum distance of the dual of $C_K$ is 2, which can be seen from the nondegenerate property of the trace function and the observation below: 

Since $\omega^{\frac{q^m-1}{q-1}} \in D$, there exist $(x_1,y_1),(x_2,y_2) \in K$ such that $x_1/x_2=y_1/y_2=\beta \in \mathbb{F}_q^*$. Then $\text{Tr}^{m_1}_1(ax_1)+\text{Tr}^m_1(by_1)=\beta \left(\text{Tr}^{m_1}_1(ax_2)+\text{Tr}^m_1(by_2) \right)$ for all $a \in \mathbb{F}_{q^{m_1}}$, $b \in \mathbb{F}_{q^m}$.

\section{Concluding remarks} \label{clu}
In this paper, inspired by the work in \cite{liu2017several} and \cite{liu2018several}, we presented a class of linear codes over $\mathbb{F}_q$ and determined their weight distributions using Gauss periods.
Our results showed that the presented linear codes have few weights and some of them are optimal or almost optimal with respect to the Griesmer bound.
It would be interesting if more linear codes with few weights can be presented.


\bibliographystyle{plain} 
\bibliography{thesis} 
\end{document}